OpenStudy (anonymous):

What amount of heat is exchanged when 16.02 grams of substance Y goes from a gas at 62 degrees Celsius to a liquid at 62 degrees Celsius? Boiling point of substance Y = 62 degrees C; ΔHvaporization = 9.22 J/mol; ΔHfusion = 2.77 J/mol. MWsubstance y = 14.77 g/mol.

OpenStudy (anonymous):

ok first calculate amount of Y 16,02/14,77 = 1,08 mol then just multiply DH fusion with that amount -> 2,77 J mol-1 * 1,08 mol =2,99 J ~ 3J

OpenStudy (anonymous):

3.00 x 10^0 J, Since substance Y is going from gas to liquid, you want the fusion value (vaporization is the opposite). But even before that, you need to convert your mass from grams to moles through the molecular weight (MW): (16.02g)/(14.77g/mol) = 1.08463 moles Now, we can find the heat by multiplying: (1.08463 moles) x (2.77 J/mol) = 3.00 x 10^0 J.